OpenStudy (anonymous):

Remainder(14! divided by 17) ??

OpenStudy (anonymous):

14! = 14*13*12*11*10*9*8*7*6*5*4*3*2*1

OpenStudy (anonymous):

help needed folks!!

hartnn (hartnn):

maybe wilson theorem might help, check that out.

mathslover (mathslover):

\[\large{\frac{14*13*12*11*10*9*8*7*6*5*4*3*2*1}{17}}\] Remainder theorem will not work here... at least

OpenStudy (anonymous):

wilson theorem shud solve dis..plz elaborate

hartnn (hartnn):

(n-1)!+1 = 0 mod n (17-1)!+1 = 0 mod 17 (16)!+1 = 0mod 17 now i m stuck.

OpenStudy (anonymous):

wot's mod n =??

mathslover (mathslover):

well my answer is "very very very long"

mathslover (mathslover):

and hopefully wrong .. :P

OpenStudy (anonymous):

please post it!!! any help wud be appreciated

mathslover (mathslover):

Remainder can not be > or equal to 17 and hence my answer is wrong @jaguarhunter007 sorry... :(

mathslover (mathslover):

Wolfram says that as *8* ..

mathslover (mathslover):

I think hartnn has a method but thats' too complicated for me now..

OpenStudy (anonymous):

Just reduce the expression.\[ 16!=16\mod17\\ 240\times14!\mod17=16\mod17\\ 14!=\frac{16}2\mod17\\ \ \ \ \ =8\mod17 \]

OpenStudy (anonymous):

oldrin where did the 2 come from?? fine till step 2..where did 240 disappear??

OpenStudy (anonymous):

\[240=2+238=2+17\times14\\240\mod17=2\]

OpenStudy (anonymous):

wow oldrin thanks!!!

OpenStudy (anonymous):

@hartnn was on the right track... \[ 16!+1=0\mod17\\ 16!=(-1)\mod17\\ \ \ \ \ \cong16\mod17 \]

OpenStudy (anonymous):

@oldrin can u give link to properties of mod...how can we move 1 to the left of equation like dat?

OpenStudy (anonymous):

oldrin u are the best!